OpenStudy (anonymous):

what is the derivative of the square root of 24x

OpenStudy (anonymous):

\[\sqrt{24x}\] can be written as \[(24x)^\frac{ 1 }{ 2}\]

OpenStudy (mertsj):

\[f'(x)=\frac{1}{2}\sqrt{24x}^{^{-\frac{1}{2}}}(24)\]

OpenStudy (anonymous):

so right, sorry doing too many questions at once, thanks @Mertsj

OpenStudy (anonymous):

\[f'(x) = \frac{ 1 }{ 2 }(24x)^\frac{ -1 }{ 2} (24)\]

OpenStudy (anonymous):

\[f'(x)= \frac{ 12 }{ \sqrt{24x} }\]